TipTop Flight School offers flying lessons at a small municipal airport. The school’s owner and manager has been attempting to evaluate performance and control costs using a variance report that compares the planning budget to actual results. A recent variance report appears below:

TipTop Flight School
Variance Report
For the Month Ended July 31
Actual
Results
Planning
Budget
Variances
Lessons 210 205
Revenue $ 54,370 $ 53,300 $ 1,070 F
Expenses:
Instructor wages 15,550 15,375 175 U
Aircraft depreciation 6,720 6,560 160 U
Fuel 4,155 3,485 670 U
Maintenance 3,705 3,530 175 U
Ground facility expenses 2,605 2,615 10 F
Administration 3,790 3,930 140 F
Total expense 36,525 35,495 1,030 U
Net operating income $ 17,845 $ 17,805 $ 40 F

After several months of using such variance reports, the owner has become frustrated. For example, she is quite confident that instructor wages were very tightly controlled in July, but the report shows an unfavorable variance.

The planning budget was developed using the following formulas, where q is the number of lessons sold:

Cost Formulas
Revenue $260q
Instructor wages $75q
Aircraft depreciation $32q
Fuel $17q
Maintenance $660 + $14q
Ground facility expenses $2,000 + $3q
Administration $3,520 + $2q

  
Required:

2. Complete the flexible budget performance report for the school for July. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

TipTop Flight School offers flying lessons at a small municipal airport. The school’s owner and manager has been attempting to evaluate performance and control costs using a variance report that compares the planning budget to actual results. A recent variance report appears below:

TipTop Flight School
Variance Report
For the Month Ended July 31
Actual
Results
Planning
Budget
Variances
Lessons 210 205
Revenue $ 54,370 $ 53,300 $ 1,070 F
Expenses:
Instructor wages 15,550 15,375 175 U
Aircraft depreciation 6,720 6,560 160 U
Fuel 4,155 3,485 670 U
Maintenance 3,705 3,530 175 U
Ground facility expenses 2,605 2,615 10 F
Administration 3,790 3,930 140 F
Total expense 36,525 35,495 1,030 U
Net operating income $ 17,845 $ 17,805 $ 40 F

After several months of using such variance reports, the owner has become frustrated. For example, she is quite confident that instructor wages were very tightly controlled in July, but the report shows an unfavorable variance.

The planning budget was developed using the following formulas, where q is the number of lessons sold:

Cost Formulas
Revenue $260q
Instructor wages $75q
Aircraft depreciation $32q
Fuel $17q
Maintenance $660 + $14q
Ground facility expenses $2,000 + $3q
Administration $3,520 + $2q

  
Required:

2. Complete the flexible budget performance report for the school for July. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

TipTop Flight School
Flexible Budget Performance Report
For the Month Ended July 31
Actual Results Flexible Budget    Planning Budget
Lessons 210    205
Revenue $54,370 $53,300
Expenses:
Instructor wages $15,550 15,375
Aircraft depreciation 6,720 6,560
Fuel 4,155 3,485
Maintenance 3,705 3,530
Ground facility expenses 2,605 2,615
Administration 3,790 3,930
Total expense 36,525 35,495
Net operating income $17,845 $17,805
Answer #1

Answer

  • Requirement

Actual results

Revenue & Spending Variance

Flexible Budget

Activity Variance

Planning Budget

Lessons

210

210

205

Revenues

$            54,370.00

$                 230.00

Unfavorable

$          54,600.00

$              1,300.00

Favourable

$              53,300.00

Expenses:

Instructer wages

$            15,550.00

$                 200.00

Favourable

$          15,750.00

$                  375.00

Unfavorable

$              15,375.00

Aircraft Depreciation

$              6,720.00

$                          -  

None

$             6,720.00

$                  160.00

Unfavorable

$                 6,560.00

Fuel

$              4,155.00

$                 585.00

Unfavorable

$             3,570.00

$                    85.00

Unfavorable

$                 3,485.00

Maintenance

$              3,705.00

$                 105.00

Unfavorable

$             3,600.00

$                    70.00

Unfavorable

$                 3,530.00

Ground Facility Expenses

$              2,605.00

$                   25.00

Favourable

$             2,630.00

$                    15.00

Unfavorable

$                 2,615.00

Administration

$              3,790.00

$                 150.00

Favourable

$             3,940.00

$                    10.00

Unfavorable

$                 3,930.00

Total Expenses

$            36,525.00

$                 315.00

Unfavorable

$          36,210.00

$                  715.00

Unfavorable

$              35,495.00

Net Operating Income

$            17,845.00

$                 545.00

Unfavorable

$          18,390.00

$                  585.00

Favourable

$              17,805.00

--Working

Actual results

Flexible Budget

Planning Budget

Lessons

210

210

205

Revenues

54370

=210*260

=205*260

Expenses:

Instructer wages

15550

=210*75

=205*75

Aircraft Depreciation

6720

=210*32

=205*32

Fuel

4155

=210*17

=205*17

Maintenance

3705

=660+(210*14)

=660+(205*14)

Ground Facility Expenses

2605

=2000+(210*3)

=2000+(205*3)

Administration

3790

=+F1887-C1887

Favourable

=3520+(210*2)

=+F1887-I1887

Unfavorable

=3520+(205*2)